# Scaling Laws for Optimal Data Mixtures

### v1

@@ Abstract
Large foundation models are typically trained on data from multiple domains, with the data mixture–the proportion of each domain used–playing a critical role in model performance. The standard approach to selecting this mixture relies on trial and error, which becomes impractical for large-scale pretraining. We propose a systematic method to determine the optimal data mixture for any target domain using scaling laws. Our approach accurately predicts the loss of a model of size $N$ trained with $D$ tokens and a specific domain weight vector $h$. We validate the universality of these scaling laws by demonstrating their predictive power in three distinct and large-scale settings: large language model (LLM), native multimodal model (NMM), and large vision models (LVM) pretraining. We further show that these scaling laws can extrapolate to new data mixtures and across scales: their parameters can be accurately estimated using a few small-scale training runs, and used to estimate the performance at larger scales and unseen domain weights. The scaling laws allow to derive the optimal domain weights for any target domain under a given training budget ($N$,$D$), providing a principled alternative to costly trial-and-error methods.
Figure 1: Scaling Laws for Optimal Data Mixtures. Left: We derive scaling laws that predict the loss of a model as a function of model size N, number of training tokens D, and the domain weights used to train the model (represented by the color of each point). The scaling law is fitted with small-scale runs with different domain weights, and used to predict accurately the loss of large-scale models trained with new, unseen domain weights. Right: We find the data mixture scaling law based on small-scale experiments (e.g., below 1B parameters) and use it to predict the optimal data mixture at larger scales (e.g., 8B parameters). Both our additive (eq. 4) and joint (eq. 5) laws lead to similar performance, and better than other mixtures (in the gray area). FLOPs are computed as 6ND.
@@ 1 Introduction
Modern machine learning models [8, 16, 20] are pre-trained on diverse data domains, such as text for LLMs, images for vision models, and mixed modalities for multimodal models. For LLMs, these domains encompass general knowledge, code, reasoning, multilingual content, and more [25, 16, 58, 5, 59]. Multimodal models [43, 69, 34, 1, 55] are trained on a mix of text, paired, and interleaved multimodal data, and finally, large vision models are trained on image domains of different qualities, containing or not images paired with text [47, 17, 20].
The domain weights determine the proportion of each domain used during training, significantly impacting model performance. However, these weights are typically chosen through ad-hoc trial and error, involving training with different domain weights and selecting what works best  [16, 55, 43]. Despite their critical role, a principled method for selecting domain weights is largely absent.
Scaling laws provide a theoretical framework to predict model performance. Initially developed for LLMs [32, 28, 30], these laws model the loss of a model as a function of the number of parameters $N$ and training tokens $D$. This framework has been extended to other domains and modalities [55, 3] and to account for factors such as the number of experts in mixture-of-experts models [33], sparsity [2], data repetitions [46], fine-tuning tokens [68, 7], and learning rate schedules [41].
@@
Scaling laws that extrapolate. We demonstrate that our scaling laws can be fitted using small-scale runs, and then provide an accurate prediction of the loss of large-scale models trained with new, unseen domain weights. This is illustrated by Fig. 1, left, where we report the loss on text when training NMMs. As expected, more text data helps reduce that loss. Our scaling laws precisely quantify this phenomenon.
Optimal domain weights estimation. Once fitted, these scaling laws give an accurate estimation of the loss as a function of the domain weights. Minimizing this estimation gives us optimal domain weights. This approach provides a principled alternative to the costly practice of trying different domain weights and selecting the best one. This is illustrated by Fig. 1, right, where we report the average loss of NMMs.
This paper is organized as follows. In Sec. 2, we introduce the problem of domain weight selection and describe our scaling law formulations. In Sec. 3, we detail the model architectures and data domains for LLM, NMM, and LVM pretraining. Fig. 3 demonstrates that our scaling laws accurately extrapolate to new domain weights, larger model sizes and number of tokens. Sec. 5 shows how the fitted laws can be used to estimate the optimal domain weights. with a few small-scale runs. Finally, Sec. 6 explores various aspects of our scaling laws, including showing that we need a small number of different domain weights to get a satisfying estimation, how the best domain weights change when scaling flops, as well as alternative scaling laws formulations. Finally, we discuss related works in Sec. 7.
@@ 2.1 Problem setup
$L_{h}(\theta)=\mathbb{E}_{x\sim\mathrm{mix}(h)}\left[\ell(x,\theta)\right]\text{ and }L_{T}(\theta)=\mathbb{E}_{x\sim\mathcal{D}_{T}}\left[\ell(x,\theta)\right]$ (1)
@@ 2.2 Scaling laws derivation
In their original form, scaling laws allow us to predict the training loss of a model of a given size $N$ after having been trained on $D$ tokens [32]. The Chinchilla scaling law models training loss as an additive power law [30]:
$\mathcal{L}(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}\enspace,$ (2)
where $E,A,\alpha$ and $\beta$ are parameters that depend on the training set, model’s architecture, and optimization algorithm. We depart from these original scaling laws in two ways: i) we consider the loss of a model on a target domain that need not be the training domain, and more importantly, ii) we quantify the impact of the domain weights $h$ on the loss. Regarding i), as already been shown in several works [45, 28, 23, 55], the loss on a target domain can still be modeled by a scaling law of the form eq. 2. Hence, for every domain weights $h$ used for training, we expect the loss on the target domain to follow a Chinchilla power law, where the coefficients depend on $h$. In other words, the loss on the target domain can be expressed as:
$\mathcal{L}(N,D,h)=E^{h}+\frac{A^{h}}{N^{\alpha^{h}}}+\frac{B^{h}}{D^{\beta^{h}}}.$ (3)
@@
$\mathcal{L}=E+\frac{1}{\sum_{i=1}^{k}C_{i}{h_{i}}^{\gamma_{i}}}+\frac{A}{{N}^{\alpha}}+\frac{B}{{D}^{\beta}}.$ (4)
@@
$\mathcal{L}=E+\frac{1}{\sum_{i=1}^{k}C_{i}h_{i}^{\gamma_{i}}}+\frac{A^{h}}{N^{\alpha}}+\frac{B^{h}}{D^{\beta}}\text{ with }A^{h}=(\sum_{i=1}^{k}C^{A}_{i}h_{i})^{\gamma^{A}}\text{ and }B^{h}=(\sum_{i=1}^{k}C_{i}^{B}h_{i})^{\gamma^{B}}$ (5)
In that scaling law, we consider the same dependency in $h$ for the bias term $E$ as in the eq. 4, and we additionally model the terms $A^{h}$ and $B^{h}$ as simple functions of $h$. The parameters of the law are $Z=(E,\alpha,\beta,(C_{i})_{i=1}^{k},(\gamma_{i})_{i=1}^{k},(C_{i}^{A})_{i=1}^{k},\gamma^{A},(C_{i}^{B})_{i=1}^{k},\gamma^{B})$, which gives $5+4k$ parameters. In this law, there is an interaction between $N$, $D$ and $h$, in the sense that $\frac{\partial^{2}\mathcal{L}}{\partial N\partial h}\neq 0$ and $\frac{\partial^{2}\mathcal{L}}{\partial D\partial h}\neq 0$, while these partial derivatives are 00 for the additive scaling law. This law predicts that the contribution of N and D to the loss depends on the domain weights, and as such, the optimal domain weights are compute-dependent. The joint scaling law is more expressive than the additive scaling law, since we can recover eq. 4 by taking $\gamma^{A}=\gamma^{B}=1$ and $C^{A}_{i}=A$, $C^{B}_{i}=B$ for all domains $i$. As such, if the scaling laws are fitted properly, the error on the training runs is always lower for the joint scaling law than for the additive scaling law. The joint scaling law still models the terms $\alpha^{h}$ and $\beta^{h}$ as constants. We tried modelling these terms using the same simple parametric form depending on $h$, but it never yielded any significant improvement (see Sec. 6). On the other hand, going from the additive to the joint law often decreased the estimation error significantly. Hence, we restrict the bulk of our study to these two laws.
@@ 2.3 Fitting the scaling laws
Figure 2: Value of the Huber loss (6) as a function of the number of L-BFGS calls to fit eq. 5 on the Interleaved domain from the multimodal experiment ($p=1062$ input-target pairs, $k=3$ domains). We repeat 100 random trials, the bold line is the median, and the shaded regions are the $25$-$75$% quantiles. The Basin-hopping method with L-BFGS subroutine converges faster than repeated calls to L-BFGS.
@@
$H(Z)=\frac{1}{p}\sum_{j=1}^{p}\mathrm{Huber}\left(L_{T}^{j}-\mathcal{L}(N^{j},D^{j},h^{j};Z)\right),$ (6)
where $\mathrm{Huber}(x)=\frac{x^{2}}{2}$ if $|x|<\delta$ and $\delta(|x|-\frac{\delta}{2})$ otherwise, where $\delta$ is a hyperparameter which we take as $\delta=1$e-$3$.
The standard technique to fit scaling laws consists of using L-BFGS [38] to minimize the loss starting from an evenly-spaced grid of initial parameters $Z$ and then retaining the smallest local minimum. Contrary to most scaling laws that involve only $5$ parameters, our scaling laws have $5+2k$ or $5+4k$ parameters, where $k$ is the number of domains. In our experiments, we use up to $k=8$ domains, which gives $37$ parameters to fit. This increased dimensionality makes the standard technique to fit scaling laws cumbersome. We propose two changes that lead to good fit. Firstly, we use a random search, to sample the initial parameters $Z$. Secondly, we use the Basin-hopping algorithm [62] instead of L-BFGS to explore the minimizers of the loss function. The Basin-hopping algorithm itself uses L-BFGS as an inner routine to minimize the loss function, but it also uses a random walk to explore the space of local minima. Fig. 2 gives an example of the performance of the algorithm: to reach a low fitting loss, the Basin-hopping algorithm requires far fewer calls to L-BFGS than doing a random search over the L-BFGS initializations.
@@ 3 Experimental setup
We give an overview of the models and domains used in our experiments. Detailed architectures and hyperparameters are given in appendix A.
@@ 3.1 Pretraining of large language models (LLMs)
Models. We use transformers [61] for autoregressive language modelling. We use the same setup as llama [60], with rotary positional embeddings, SwiGLU activations, and RMSNorm. The models are scaled by changing the latent dimension, with model sizes ranging from 186M to 7B parameters.
For some smaller-scale analyses, we also use GPT2-style transformers [50] to perform autoregressive language modeling with model sizes ranging from 90M to 3B parameters.
Datasets. For the main experiments, we use the $k=7$ domains from slimpajama [56]. We use these domains as distributed by the authors, without any additional data filtering.
For some smaller-scale analyses, we use up to $k=8$ domains coming from the Pile dataset [21]: Wikipedia, StackExchange, GitHub, pg19, arxiv, free law, openwebtext, and PubMed Central.
@@ 3.2 Pretraining of native multimodal models (NMMs)
Models. We pretrain native multimodal models (NMMs), based on an early-fusion architecture [6, 55] and follow the design and implementation proposed in [55]. The model consists of a single transformer [61] without a separate vision encoder, resulting in the same architecture used for LLMs. The model processes a sequence of interleaved text and image tokens. Text tokens are obtained using a standard LLM tokenizer, while image tokens are obtained by patchifying the image and applying a linear projection. Images are resized to 224×224 resolution with a 14×14 patch size. The overall model architecture is aligned with [35], incorporating SwiGLU FFNs [53] and QK-Norm [15].
Datasets. Following previous works [55, 34, 37] we train on a mixture of multimodal datasets, covering $k=3$ data types: (1) text-only data from DCLM [35], (2) interleaved multimodal documents from Obelics [34], and (3) paired image-caption datasets from DFN [19], COYO [10], and a private collection of High-Quality Image-Text Pairs (HQITP).
@@ 3.3 Pretraining of large vision models (LVMs)
Models. We pretrain large vision models with a multimodal objective, following the AIMv2 recipe [20]. Unlike traditional language modeling or multimodal models described above that focus on text decoding, AIMv2 trains a vision encoder using an autoregressive objective on both image and text tokens. The model architecture is composed of a vision encoder and a multimodal decoder stitched together in a late-fusion fashion.
Datasets. We train on a mixture of paired of image-caption datasets drawn from four domains ($k=4$): (1) noisy alt-text sourced from the Internet, including COYO-700M [10] and DFN2B [19], which provide large-scale real-world image-text pairs with varying levels of noise and quality; (2) HQ-ITP-1, a high-quality dataset containing 134 million samples; (3) HQ-ITP-2, another high-quality dataset comprising 400 million samples; and (4) synthetic data, consisting of recaptioned versions of DFN2B and HQ-ITP-2.

### v2

@@ Abstract
Large foundation models are typically trained on data from multiple domains, with the data mixture–the proportion of each domain used–playing a critical role in model performance. The standard approach to selecting this mixture relies on trial and error, which becomes impractical for large-scale pretraining. We propose a systematic method to determine the optimal data mixture for any target domain using scaling laws. Our approach accurately predicts the loss of a model of size $N$ trained with $D$ tokens and a specific domain weight vector $h$. We validate the universality of these scaling laws by demonstrating their predictive power in three distinct and large-scale settings: large language model (LLM), native multimodal model (NMM), and large vision models (LVM) pretraining. We further show that these scaling laws can extrapolate to new data mixtures and across scales: their parameters can be accurately estimated using a few small-scale training runs, and used to estimate the performance at larger scales and unseen domain weights. The scaling laws allow to derive the optimal domain weights for any target domain under a given training budget (NN,$D$), providing a principled alternative to costly trial-and-error methods.
[…]
[Correspondence]Pierre Ablin: p_ablin@apple.com
Apple and the Apple logo are trademarks of Apple Inc., registered in the U.S. and other countries and regions.
Refer to captionRefer to caption
Figure 1: Scaling Laws for Optimal Data Mixtures. Left: We derive scaling laws that predict the loss of a model as a function of model size N, number of training tokens D, and the domain weights used to train the model (represented by the color of each point). The scaling law is fitted with small-scale runs with different domain weights, and used to predict accurately the loss of large-scale models trained with new, unseen domain weights. Right: We find the data mixture scaling law based on small-scale experiments (e.g., below 1B parameters) and use it to predict the optimal data mixture at larger scales (e.g., 8B parameters). Both our additive (equation˜2.4) and joint (equation˜2.5) laws lead to similar performance, and better than other mixtures (in the gray area). FLOPs are computed as 6ND.
@@ 1 Introduction
Modern machine learning models (Brown et al., 2020; Dubey et al., 2024; Fini et al., 2024) are pre-trained on diverse data domains, such as text for LLMs, images for vision models, and mixed modalities for multimodal models. For LLMs, these domains encompass general knowledge, code, reasoning, multilingual content, and more (Grangier et al., 2024; Dubey et al., 2024; Team et al., 2023; Bai et al., 2023; Team et al., 2024). Multimodal models (McKinzie et al., 2024; Zhang et al., 2024b; Laurençon et al., 2024; Abdin et al., 2024; Shukor et al., 2025) are trained on a mix of text, paired, and interleaved multimodal data, and finally, large vision models are trained on image domains of different qualities, containing or not images paired with text (Oquab et al., 2023; El-Nouby et al., 2024; Fini et al., 2024).
The domain weights determine the proportion of each domain used during training, significantly impacting model performance. However, these weights are typically chosen through ad-hoc trial and error, involving training with different domain weights and selecting what works best (Dubey et al., 2024; Shukor et al., 2025; McKinzie et al., 2024). Despite their critical role, a principled method for selecting domain weights is largely absent.
Scaling laws provide a theoretical framework to predict model performance. Initially developed for LLMs (Kaplan et al., 2020; Hernandez et al., 2021; Hoffmann et al., 2022), these laws model the loss of a model as a function of the number of parameters $N$ and training tokens $D$. This framework has been extended to other domains and modalities (Shukor et al., 2025; Aghajanyan et al., 2023) and to account for factors such as the number of experts in mixture-of-experts models (Krajewski et al., 2024), sparsity (Abnar et al., 2025), data repetitions (Muennighoff et al., 2023), fine-tuning tokens (Zhang et al., 2024a; Bethune et al., 2025), and learning rate schedules (Luo et al., 2025).
@@
Scaling laws that extrapolate. We demonstrate that our scaling laws can be fitted using small-scale runs, and then provide an accurate prediction of the loss of large-scale models trained with new, unseen domain weights. This is illustrated by figure˜1, left, where we report the loss on text when training NMMs. As expected, more text data helps reduce that loss. Our scaling laws precisely quantify this phenomenon.
Optimal domain weights estimation. Once fitted, these scaling laws give an accurate estimation of the loss as a function of the domain weights. Minimizing this estimation gives us optimal domain weights. This approach provides a principled alternative to the costly practice of trying different domain weights and selecting the best one. This is illustrated by figure˜1, right, where we report the average loss of NMMs.
This paper is organized as follows. In section˜2, we introduce the problem of domain weight selection and describe our scaling law formulations. In section˜3, we detail the model architectures and data domains for LLM, NMM, and LVM pretraining. figure˜3 demonstrates that our scaling laws accurately extrapolate to new domain weights, larger model sizes and number of tokens. section˜5 shows how the fitted laws can be used to estimate the optimal domain weights. with a few small-scale runs. Finally, section˜6 explores various aspects of our scaling laws, including showing that we need a small number of different domain weights to get a satisfying estimation, how the best domain weights change when scaling flops, as well as alternative scaling laws formulations. Finally, we discuss related works in section˜7.
@@ 2.1 Problem setup
$L_{h}(\theta)=\mathbb{E}_{x\sim\mathrm{mix}(h)}\left[\ell(x,\theta)\right]\text{ and }L_{T}(\theta)=\mathbb{E}_{x\sim\mathcal{D}_{T}}\left[\ell(x,\theta)\right]$ (2.1)
@@ 2.2 Scaling laws derivation
In their original form, scaling laws allow us to predict the training loss of a model of a given size $N$ after having been trained on $D$ tokens (Kaplan et al., 2020). The Chinchilla scaling law models training loss as an additive power law (Hoffmann et al., 2022):
$\mathcal{L}(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}\kern 5.0pt,$ (2.2)
where $E,A,\alpha$ and $\beta$ are parameters that depend on the training set, model’s architecture, and optimization algorithm. We depart from these original scaling laws in two ways: i) we consider the loss of a model on a target domain that need not be the training domain, and more importantly, ii) we quantify the impact of the domain weights $h$ on the loss. Regarding i), as already been shown in several works (Mikami et al., 2022; Hernandez et al., 2021; Ghorbani et al., 2021; Shukor et al., 2025), the loss on a target domain can still be modeled by a scaling law of the form equation˜2.2. Hence, for every domain weights $h$ used for training, we expect the loss on the target domain to follow a Chinchilla power law, where the coefficients depend on $h$. In other words, the loss on the target domain can be expressed as:
$\mathcal{L}(N,D,h)=E^{h}+\frac{A^{h}}{N^{\alpha^{h}}}+\frac{B^{h}}{D^{\beta^{h}}}.$ (2.3)
@@
$\mathcal{L}=E+\frac{1}{\sum_{i=1}^{k}C_{i}{h_{i}}^{\gamma_{i}}}+\frac{A}{{N}^{\alpha}}+\frac{B}{{D}^{\beta}}.$ (2.4)
@@
$\mathcal{L}=E+\frac{1}{\sum_{i=1}^{k}C_{i}h_{i}^{\gamma_{i}}}+\frac{A^{h}}{N^{\alpha}}+\frac{B^{h}}{D^{\beta}}\text{ with }A^{h}=(\sum_{i=1}^{k}C^{A}_{i}h_{i})^{\gamma^{A}}\text{ and }B^{h}=(\sum_{i=1}^{k}C_{i}^{B}h_{i})^{\gamma^{B}}$ (2.5)
In that scaling law, we consider the same dependency in $h$ for the bias term $E$ as in the equation˜2.4, and we additionally model the terms $A^{h}$ and $B^{h}$ as simple functions of $h$. The parameters of the law are $Z=(E,\alpha,\beta,(C_{i})_{i=1}^{k},(\gamma_{i})_{i=1}^{k},(C_{i}^{A})_{i=1}^{k},\gamma^{A},(C_{i}^{B})_{i=1}^{k},\gamma^{B})$, which gives $5+4k$ parameters. In this law, there is an interaction between $N$, $D$ and $h$, in the sense that $\frac{\partial^{2}\mathcal{L}}{\partial N\partial h}\neq 0$ and $\frac{\partial^{2}\mathcal{L}}{\partial D\partial h}\neq 0$, while these partial derivatives are 0 for the additive scaling law. This law predicts that the contribution of N and D to the loss depends on the domain weights, and as such, the optimal domain weights are compute-dependent. The joint scaling law is more expressive than the additive scaling law, since we can recover equation˜2.4 by taking $\gamma^{A}=\gamma^{B}=1$ and $C^{A}_{i}=A$, $C^{B}_{i}=B$ for all domains $i$. As such, if the scaling laws are fitted properly, the error on the training runs is always lower for the joint scaling law than for the additive scaling law. The joint scaling law still models the terms $\alpha^{h}$ and $\beta^{h}$ as constants. We tried modelling these terms using the same simple parametric form depending on $h$, but it never yielded any significant improvement (see section˜6). On the other hand, going from the additive to the joint law often decreased the estimation error significantly. Hence, we restrict the bulk of our study to these two laws.
@@ 2.3 Fitting the scaling laws
Figure 2: Value of the Huber loss (2.6) as a function of the number of L-BFGS calls to fit equation˜2.5 on the Interleaved domain from the multimodal experiment ($p=1062$ input-target pairs, $k=3$ domains). We repeat 100 random trials, the bold line is the median, and the shaded regions are the 2525-7575% quantiles. The Basin-hopping method with L-BFGS subroutine converges faster than repeated calls to L-BFGS.
@@
$H(Z)=\frac{1}{p}\sum_{j=1}^{p}\mathrm{Huber}\left(L_{T}^{j}-\mathcal{L}(N^{j},D^{j},h^{j};Z)\right),$ (2.6)
where $\mathrm{Huber}(x)=\frac{x^{2}}{2}$ if $|x|<\delta$ and $\delta(|x|-\frac{\delta}{2})$ otherwise, where $\delta$ is a hyperparameter which we take as $\delta=1$e-33.
The standard technique to fit scaling laws consists of using L-BFGS (Liu and Nocedal, 1989) to minimize the loss starting from an evenly-spaced grid of initial parameters $Z$ and then retaining the smallest local minimum. Contrary to most scaling laws that involve only $5$ parameters, our scaling laws have $5+2k$ or $5+4k$ parameters, where $k$ is the number of domains. In our experiments, we use up to $k=8$ domains, which gives $37$ parameters to fit. This increased dimensionality makes the standard technique to fit scaling laws cumbersome. We propose two changes that lead to good fit. Firstly, we use a random search, to sample the initial parameters $Z$. Secondly, we use the Basin-hopping algorithm (Wales and Doye, 1997) instead of L-BFGS to explore the minimizers of the loss function. The Basin-hopping algorithm itself uses L-BFGS as an inner routine to minimize the loss function, but it also uses a random walk to explore the space of local minima. figure˜2 gives an example of the performance of the algorithm: to reach a low fitting loss, the Basin-hopping algorithm requires far fewer calls to L-BFGS than doing a random search over the L-BFGS initializations.
@@ 3 Experimental setup
We give an overview of the models and domains used in our experiments. Detailed architectures and hyperparameters are given in appendix˜A.
@@ 3.1 Pretraining of large language models (LLMs)
Models. We use transformers (Vaswani, 2017) for autoregressive language modelling. We use the same setup as llama (Touvron et al., 2023), with rotary positional embeddings, SwiGLU activations, and RMSNorm. The models are scaled by changing the latent dimension, with model sizes ranging from 186M to 7B parameters.
For some smaller-scale analyses, we also use GPT2-style transformers (Radford et al., 2019) to perform autoregressive language modeling with model sizes ranging from 90M to 3B parameters.
Datasets. For the main experiments, we use the $k=7$ domains from slimpajama (Soboleva et al., 2023). We use these domains as distributed by the authors, without any additional data filtering.
For some smaller-scale analyses, we use up to $k=8$ domains coming from the Pile dataset (Gao et al., 2020): Wikipedia, StackExchange, GitHub, pg19, arxiv, free law, openwebtext, and PubMed Central.
@@ 3.2 Pretraining of native multimodal models (NMMs)
Models. We pretrain native multimodal models (NMMs), based on an early-fusion architecture (Bavishi et al., 2023; Shukor et al., 2025) and follow the design and implementation proposed in (Shukor et al., 2025). The model consists of a single transformer (Vaswani, 2017) without a separate vision encoder, resulting in the same architecture used for LLMs. The model processes a sequence of interleaved text and image tokens. Text tokens are obtained using a standard LLM tokenizer, while image tokens are obtained by patchifying the image and applying a linear projection. Images are resized to 224×224 resolution with a 14×14 patch size. The overall model architecture is aligned with (Li et al., 2024), incorporating SwiGLU FFNs (Shazeer, 2020) and QK-Norm (Dehghani et al., 2023).
Datasets. Following previous works (Shukor et al., 2025; Laurençon et al., 2024; Lin et al., 2024) we train on a mixture of multimodal datasets, covering $k=3$ data types: (1) text-only data from DCLM (Li et al., 2024), (2) interleaved multimodal documents from Obelics (Laurençon et al., 2024), and (3) paired image-caption datasets from DFN (Fang et al., 2023), COYO (Byeon et al., 2022), and a private collection of High-Quality Image-Text Pairs (HQITP).
@@ 3.3 Pretraining of large vision models (LVMs)
Models. We pretrain large vision models with a multimodal objective, following the AIMv2 recipe (Fini et al., 2024). Unlike traditional language modeling or multimodal models described above that focus on text decoding, AIMv2 trains a vision encoder using an autoregressive objective on both image and text tokens. The model architecture is composed of a vision encoder and a multimodal decoder stitched together in a late-fusion fashion.
Datasets. We train on a mixture of paired of image-caption datasets drawn from four domains ($k=4$): (1) noisy alt-text sourced from the Internet, including COYO-700M (Byeon et al., 2022) and DFN2B (Fang et al., 2023), which provide large-scale real-world image-text pairs with varying levels of noise and quality; (2) HQ-ITP-1, a high-quality dataset containing 134 million samples; (3) HQ-ITP-2, another high-quality dataset comprising 400 million samples; and (4) synthetic data, consisting of recaptioned versions of DFN2B and HQ-ITP-2.